# Neural Risk Limiting Dispatch in Power Networks: Formulation and Generalization Guarantees

### v1

@@ Abstract
Risk limiting dispatch (RLD) has been proposed as an approach that effectively trades off economic costs with operational risks for power dispatch under uncertainty. However, how to solve the RLD problem with provably near-optimal performance still remains an open problem. This paper presents a learning-based solution to this challenge. We first design a data-driven formulation for the RLD problem, which aims to construct a decision rule that directly maps day-ahead observable information to cost-effective dispatch decisions for the future delivery interval. Unlike most existing works that follow a predict-then-optimize paradigm, this end-to-end rule bypasses the additional suboptimality introduced by separately handling prediction and optimization. We then propose neural RLD, a novel solution method to the data-driven formulation. This method leverages an L2-regularized neural network to learn the decision rule, thereby transforming the data-driven formulation into a neural network training task that can be efficiently completed by stochastic gradient descent. A theoretical performance guarantee is further established to bound the suboptimality of our method, which implies that its suboptimality approaches to zero with high probability as more samples are utilized. Simulation tests across various systems demonstrate our method’s superior performance in convergence, suboptimality, and computational efficiency compared with benchmarks.
@@ I Introduction
Within the second category, Risk Limiting Dispatch (RLD) distinguishes itself as a class of methods that effectively trade off the economic cost with operation risks utilizing the most updated information about the uncertainties [4, 5]. Varaiya et al. [4] initially proposed RLD within a multi-stage stochastic control formulation to incorporate recourse decisions in multi-settlement markets. This approach was later generalized to explicitly leveraging forecast updates (i.e., the progressively improving forecasts for loads and renewable generation as we approach the delivery time) [5], generator ramping constraints [6], and energy storage [7]. These early studies mostly relied on deriving on analytical solutions for the underlying stochastic control problems, circumventing the curse of dimensionality associated with continuous state/action/disturbance variables in the original RLD formulation. However, when power network constraints are considered (e.g., as in the network RLD problem), multi-stage stochastic control formulation for RLD turns out to be intractable due to the high-dimensional continuous state space and the lack of analytical solutions. In fact, even the two-stage special case of the network RLD problem is challenging, for which Zhang et al. [8] proposed a heuristic algorithm and established its optimality under a “small-$\sigma$ assumption”, where $\sigma$ is the standard deviation of the forecast errors. It is noteworthy that the formulation and proposed algorithm in [8] do not rely on sampling from the probability distribution of the disturbances, which can be computationally costly when the disturbances are high-dimensional (e.g., considering correlated spatially disbursed wind and solar outputs). This makes their proposed algorithm easy to be incorporated in existing power system dispatch software (e.g., by modifying the reserve margin based on RLD solutions). Variants of RLD formulations have subsequently found a number of other applications in power system operations and planning [9, 10, 11]. Despite these prior studies, it remains an open problem to develop provably near-optimal methods that solve the RLD problem with network constraints and practical uncertainty levels (i.e., without the small-$\sigma$ assumption).
@@
Even if the difficulties associated with characterizing the uncertainties, i.e., the prediction (albeit probabilistic) step, and identifying the optimal decisions given the uncertainties, i.e., the optimization step, can be resolved, fundamental pitfalls around the two-step predict-then-optimize paradigm make it challenging to achieve direct data-driven, end-to-end optimal or provably near optimal RLD decisions. Indeed, it is common for the prediction step to use standard statistical loss functions, e.g., mean squared error (MSE), to quantify the accuracy ignoring the actual objective functions in the subsequent optimization step, and the optimization step to assumes the uncertainty model obtained from the prediction step is precise. However, when the errors in the prediction step is accounted for, a “more accurate” model for uncertainty as measured by MSE may not always lead to a better decision in the optimization step [17, 18]. Because of this “mismatch” in the objectives of the two steps, this paradigm may introduce additional suboptimality, resulting in undesirable performance.
@@ I-A Contributions and paper organization
In this paper, we propose neural RLD, a formulation and method that merge the prediction and optimization steps, which directly use the data to train a neural network outputting dispatch decisions. This enables us to circumvent the challenges associated with the individual steps in the overall pitfalls of the predict-then-optimize paradigm, while obtaining end-to-end generalization bounds for the quality of the resulting dispatch decisions. This paper contributes to the literature in the following ways:
@@
    We propose the neural RLD method to solve the data-driven formulation above. This method utilizes an L2-regularized neural network to learn the decision rule, effectively converting the data-driven formulation into a neural network training task with a specialized training loss. We then explicitly derive the gradient of this specialized loss, enabling efficient completion of this training task by using stochastic gradient descent (SGD).
@@
    We develop a generalization performance guarantee based on statistical learning tools to theoretically bound the suboptimality of the neural RLD method. This guarantee implies that our method is asymptotically optimal, i.e., its suboptimality is likely to gradually decrease to zero with high probability as more samples are used.
@@ I-B1 Learning-based stochastic dispatch
The application of machine learning to stochastic dispatch presents both innovative solutions and challenges. Initially, machine learning methods were usually deployed in one of two steps in the predict-then-optimize paradigm: enhancing prediction accuracy in the prediction step [19], or improving computational efficiency in the optimization step [20]. However, the aforementioned mismatch issue may bring additional suboptimality. In addressing this limitation, end-to-end frameworks, such as imitation learning [21], emerge as an alternative. By directly mapping observed features to dispatch decisions, imitation learning bypasses the mismatch issue. However, these end-to-end methods have been limited to deterministic problems [22]. Moreover, they often use the mean square error (MSE) between actual and predicted optimal decisions as the training loss. This choice might not accurately reflect the suboptimality of predicted decisions [17]. Additionally, despite demonstrating practical effectiveness in numerical case studies, these methods frequently fall short in providing theoretical performance assurances under varying conditions.
@@ I-B2 Sample average approximation
Sample average approximation (SAA) is a method [23] that uses a finite number of discrete scenarios to approximate uncertainties, which can effectively transform RLD into a deterministic form. Due to its ease of use, it has been applied to various dispatch tasks, including economic dispatch [24], unit commitment [25], and optimal power flow [26]. Meanwhile, some error bounds have also been proposed to explain its performance on characterizing uncertainty. For example, Shapiro et al. [27] have shown that the SAA solution uniformly converges to the optimal one as the number of scenarios increases indefinitely. Additionally, Kim et al. [28] established a link between the number of scenarios used and the worst-case optimality gap of the SAA. However, the SAA still suffers from the aforementioned “mismatch” issue, as it is typically incorporated in the optimization step of the predict-then-optimize paradigm.
@@ I-B3 Direct learning
In conference paper [12], we initially proposed an end-to-end learning method to predict RLD decisions directly from historical data. This method offered a generalization guarantee to bound its suboptimality. However, the initial study was limited to scalar decision variables, which significantly restricted its applicability and practicality in scenarios involving power network constraints. Meanwhile, it focused on theoretical analysis without numerical case studies. The current paper significantly extends the scope of [12] by concentrating on the network RLD with vectorial variables, providing a new generalization performance guarantee for this networked case, and conducting a series of case studies to validate the practical applicability of our method.
@@ II Problem Formulation
We start by summarizing the classical network RLD formulation pertaining to the optimization step in the predict-then-optimize paradigm. We will then introduce a data-driven formulation of the problem that merges the prediction and optimization steps.
@@ II-A Classic network risk limiting dispatch
Consider a power network with ${N_{\mathrm{B}}}$ buses. In the day-ahead stage, the SO dispatches generators based on the most updated information regarding the net demand (i.e., load and renewable generation) over the network. Denote this dispatch decision by $\mathbf{u}\in\mathbb{R}^{N_{\mathrm{B}}}$, where $u_{n}$ is the generation at bus $n$. As the actual net demand $\mathbf{d}\in\mathbb{R}^{N_{\mathrm{B}}}$ is not available at day ahead, classical RLD formulation assumes that the probability distribution of $\mathbf{d}$ can be estimated. This probability distribution can depend on various auxiliary factors (e.g., temperature and wind speed information from the day-ahead weather forecast), and thus is denoted by $\mathcal{D}(\mathbf{x})$, where $\mathbf{x}\in\mathbb{R}^{p}$ is the $p\times 1$ vector of auxiliary information available at day ahead. In the real-time stage, the SO observes the realization of net demand $\mathbf{d}$ and can adjust the day-ahead dispatch by either dispatching extra generation or reducing the day-ahead dispatched amount, so the adjusted total generation profile balances the grid. We denote the real-time adjustment by $\mathbf{g}\in\mathbb{R}^{N_{\mathrm{B}}}$. Mathematically, the two-stage network RLD problem takes the form of
@@
where $\bm{\alpha}\in\mathbb{R}^{N_{\mathrm{B}}}$ is the vector of linear generation cost coefficient for the day-ahead stage, and $Q(\mathbf{u},\mathbf{d})$ is the optimal second-stage cost given the day-ahead dispatch $\mathbf{u}$ and the realized net demand $\mathbf{d}$, defined as the optimal value of the real-time dispatch problem. In particular, as in [8], we use the following optimal DC (i.e., linearized AC) power flow formulation for the real-time dispatch:
@@
$-\mathbf{f}^{\mathrm{max}}\leq\mathbf{F}\bm{\theta}\leq\mathbf{f}^{\mathrm{max}}:\bm{\nu}_{-},\bm{\nu}_{+},$ (2c)
where the objective function models the cost of dispatching extra real-time generation or shedding loads with $(\mathbf{g})_{+}:=\max\{\mathbf{g},\mathbf{0}\}$ and $\bm{\beta}\in\mathbb{R}^{N_{\mathrm{B}}}$ modeling the per unit cost of extra generation or value of loss load. Eq. (2b) represents the DC power flow equation, where matrix $\mathbf{B}\in\mathbb{R}^{{N_{\mathrm{B}}}\times{N_{\mathrm{B}}}}$ maps voltage phase angles $\bm{\theta}\in\mathbb{R}^{N_{\mathrm{B}}}$ into nodal active power injections. Eq. (2c) denotes the branch capacity constraints, where $\mathbf{F}\in\mathbb{R}^{{N_{\mathrm{L}}}\times{N_{\mathrm{B}}}}$ maps $\bm{\theta}$ to branch flows, $\mathbf{f}^{\mathrm{max}}$ is the corresponding maximum allowable values, and ${N_{\mathrm{L}}}$ is the number of branches. Vectors $\bm{\lambda}\in\mathbb{R}^{N_{\mathrm{B}}}$, $\bm{\nu}_{-}\in\mathbb{R}^{N_{\mathrm{L}}}$ and $\bm{\nu}_{+}\in\mathbb{R}^{N_{\mathrm{L}}}$ are the dual variables of (2b)-(2c), respectively.
@@
II-B Data-driven risk limiting dispatch
@@
Ideally, we wish to find the optimal decision rule $\mathbf{h}^{\star}:\mathbb{R}^{p}\mapsto\mathbb{R}^{N_{\mathrm{B}}}$, that maps the day-ahead auxiliary information $\mathbf{x}$ to the optimal dispatch decision $\mathbf{u}^{\star}$. [^1] Given $\mathbf{u}$ and the realization of $\mathbf{d}$, it is straightforward to calculate the optimal real-time dispatch by solving (II-A). Thus it suffices to focus on the day-ahead decision $\mathbf{u}$. Adapting terminologies from the statistical learning theory, we will also refer to such a decision rule as a hypothesis. Mathematically, supposing that there is an underlying true joint distribution of $(\mathbf{x},\mathbf{d})$, denoted by $\mathcal{Z}$, we can write the following problem to identify the optimal hypothesis as

### v2

@@ Abstract
Risk limiting dispatch (RLD) has been proposed as an approach that effectively trades off economic costs with operational risks for power dispatch under uncertainty. However, how to solve the RLD problem with provably near-optimal performance still remains an open problem. This paper presents a learning-based solution to this challenge. We first design a data-driven formulation for the RLD problem, which aims to construct a decision rule that directly maps day-ahead observable information to cost-effective dispatch decisions for the future delivery interval. Unlike most existing works that follow a predict-then-optimize paradigm, this end-to-end rule bypasses the additional suboptimality introduced by separately handling prediction and optimization. We then propose neural RLD, a novel solution method to the data-driven formulation. This method leverages an L2-regularized neural network (NN) to learn the decision rule, thereby transforming the data-driven formulation into a training task that can be efficiently completed by stochastic gradient descent. A theoretical performance guarantee is further established to characterize the suboptimality of our method, which implies that its suboptimality approaches zero with high probability as more samples are utilized. Simulation tests across various systems demonstrate our method’s superior performance in convergence, suboptimality, and computational efficiency compared with benchmarks.
@@ I Introduction
Within the second category, Risk Limiting Dispatch (RLD) distinguishes itself as a class of methods that effectively trade off the economic cost with operation risks utilizing the most updated information about the uncertainties [4, 5]. Varaiya et al. [4] initially proposed RLD within a multi-stage stochastic control formulation to incorporate recourse decisions in multi-settlement markets. This approach was later generalized to explicitly leveraging forecast updates (i.e., the progressively improving forecasts for loads and renewable generation as we approach the delivery time) [5], generator ramping constraints [6], and energy storage [7]. These early studies mostly relied on deriving analytical solutions for the underlying stochastic control problems, circumventing the curse of dimensionality associated with continuous state/action/disturbance variables in the original RLD formulation. However, when power network constraints are considered (e.g., as in the network RLD problem), multi-stage stochastic control formulation for RLD turns out to be intractable due to the high-dimensional continuous state space and the lack of analytical solutions. In fact, even the two-stage special case of the network RLD problem is challenging, for which Zhang et al. [8] proposed a heuristic algorithm and established a performance guarantee to characterize its suboptimality under a “small-$\sigma$ assumption”, where $\sigma$ represents the standard deviation of forecast errors. This assumption implies that $\sigma$ is much smaller than the expected net demand. Under this condition, forecast errors have minimal impact on the network’s operational characteristics, and the dispatch decisions made under both deterministic and stochastic settings remain similar. It is noteworthy that the formulation and proposed algorithm in [8] do not rely on sampling from the probability distribution of the disturbances, which can be computationally costly when the disturbances are high-dimensional (e.g., considering correlated spatially dispersed wind and solar outputs). This makes their proposed algorithm easy to be incorporated in existing power system dispatch software (e.g., by modifying the reserve margin based on RLD solutions). However, this assumption may not hold in practice, particularly in power systems with high renewable penetration, where forecast errors can be significant. In such cases, the performance guarantee established under the small-$\sigma$ assumption may no longer apply, especially when renewable generation results in negative net demand, as the authors of [8] acknowledge. Hence, it remains an open problem to develop provably near-optimal methods that solve the RLD problem with network constraints and practical uncertainty levels (i.e., without the small-$\sigma$ assumption).
@@
Even if the difficulties associated with characterizing the uncertainties, i.e., the prediction (albeit probabilistic) step, and identifying the optimal decisions given the uncertainties, i.e., the optimization step, can be resolved, fundamental pitfalls around the two-step predict-then-optimize paradigm make it challenging to achieve direct data-driven, end-to-end optimal or provably near optimal RLD decisions. Indeed, it is common for the prediction step to use standard statistical loss functions, e.g., mean squared error (MSE), to quantify the accuracy ignoring the actual objective functions in the subsequent optimization step, and the optimization step assumes the uncertainty model obtained from the prediction step is precise. However, when the errors in the prediction step are accounted for, a “more accurate” model for uncertainty as measured by MSE may not always lead to a better decision in the optimization step [14, 15]. Because of this “mismatch” in the objectives of the two steps, this paradigm may introduce additional suboptimality, resulting in undesirable performance.
@@ I-A Contributions and paper organization
In this paper, we propose neural RLD, a formulation and method that merges the prediction and optimization steps, which directly use the data to train a neural network (NN) outputting dispatch decisions. This enables us to circumvent the challenges associated with the individual steps in the overall pitfalls of the predict-then-optimize paradigm, while obtaining end-to-end generalization bounds for the quality of the resulting dispatch decisions. This paper contributes to the literature in the following ways:
@@
    We propose the neural RLD method to solve the data-driven formulation above. This method utilizes an L2-regularized NN to learn the decision rule, effectively converting the data-driven formulation into a NN training task with a specialized training loss. We then explicitly derive the gradient of this specialized loss, enabling efficient completion of this training task by using stochastic gradient descent (SGD).
@@
    We establish a performance guarantee that theoretically characterizes the generalization error of the proposed method. This guarantee demonstrates that our method is asymptotically optimal if the NN is well-trained, implying that its optimality gap approaches zero with high probability as the sample size increases. To the best of our knowledge, this is the first performance guarantee developed for an end-to-end RLD approach.
@@ I-B1 Learning-based stochastic dispatch
The application of machine learning to stochastic dispatch presents both innovative solutions and challenges. Initially, machine learning methods were usually deployed in one of two steps in the predict-then-optimize paradigm: enhancing prediction accuracy in the prediction step [16], or improving computational efficiency in the optimization step [17]. However, the aforementioned mismatch issue may bring additional suboptimality. To address this limitation, end-to-end frameworks have emerged as a promising alternative. For example, imitation learning trains NNs to directly map observed features to dispatch decisions, thereby avoiding the mismatch issue. However, this framework typically uses mean squared error (MSE) between predicted and actual optimal decisions as the training loss, which may not effectively capture the suboptimality of the predicted decisions [14]. Some studies have instead used the original objective of the power dispatch problem as the training loss, which can directly measure suboptimality [18, 19]. Nevertheless, most of these studies are confined to deterministic power dispatch tasks. A few studies have developed learning-based end-to-end methods for stochastic problems [20, 21], relying on limited historical data to approximate the true distribution of uncertain operation conditions in future delivery intervals. This discrepancy between the empirical distribution derived from data and the true population distribution introduces a generalization error that can induce suboptimality. Yet, none of these studies has provided a theoretical performance guarantee to quantify this generalization error and assess its potential impact on suboptimality, thereby limiting their practical applicability.
@@ I-B2 Sample average approximation
Sample average approximation (SAA) is a method [22] that uses a finite number of discrete scenarios to approximate uncertainties, which can effectively transform RLD into a deterministic form. Some error bounds have also been proposed to explain its performance on characterizing uncertainty. For example, Shapiro et al. [23] have shown that the SAA solution uniformly converges to the optimal one as the number of scenarios increases indefinitely. Additionally, Kim et al. [24] established a link between the number of scenarios used and the worst-case optimality gap of the SAA. However, the SAA still suffers from the aforementioned “mismatch” issue, as it is typically incorporated in the optimization step of the predict-then-optimize paradigm.
@@ I-B3 Direct learning
In conference paper [9], we initially proposed an end-to-end learning method to predict RLD decisions directly from historical data. This method offered a generalization guarantee to characterize its suboptimality. However, the initial study was limited to scalar decision variables, which significantly restricted its applicability and practicality in scenarios involving power network constraints. Meanwhile, it focused on theoretical analysis without numerical case studies. The current paper significantly extends the scope of [9] by concentrating on the network RLD with vector variables, providing a new generalization performance guarantee for this networked case, and conducting a series of case studies to validate the practical applicability of our method.
@@ II Problem Formulation
This paper proposes a neural RLD approach, where a NN is trained to learn a decision rule that maps observable day-ahead information to cost-effective generation decisions. The detailed procedure for deriving the neural RLD formulation is summarized in Fig. 1. We start by presenting the classic network RLD (1) and the data-driven formulation (4) derived using empirical risk minimization (ERM). Then, the data-driven formulation is translated into a NN training task, resulting in the proposed neural RLD (7). The gradient of the training loss is explicitly derived using the envelope theorem, enabling efficient NN training. Furthermore, a generalization performance guarantee is established using statistical learning theory, which theoretically characterizes the suboptimality of our neural RLD.
This section introduces the classic and data-driven RLD formulations. The remaining methodology will be detailed in Section III.
Figure 1: Steps for deriving the proposed neural RLD formulation.
@@ II-A Classic network risk limiting dispatch
Consider a power network with ${N_{\mathrm{B}}}$ buses. In the day-ahead stage, the SO dispatches generators based on the most updated information regarding the net demand (i.e., load and renewable generation) over the network. Denote this dispatch decision by $\mathbf{u}\in\mathbb{R}^{N_{\mathrm{B}}}$, where $u_{n}$ is the generation at bus $n$. As the actual net demand $\mathbf{d}\in\mathbb{R}^{N_{\mathrm{B}}}$ is not available at day ahead, the classical RLD formulation assumes that the probability distribution of $\mathbf{d}$ can be estimated. This probability distribution can depend on various auxiliary factors (e.g., temperature and wind speed information from the day-ahead weather forecast), and thus is denoted by $\mathcal{D}(\mathbf{x})$, where $\mathbf{x}\in\mathbb{R}^{p}$ is the $p\times 1$ vector of auxiliary information available at day ahead. In the real-time stage, the SO observes the realization of net demand $\mathbf{d}$ and can adjust the day-ahead dispatch by either dispatching extra generation or reducing the day-ahead dispatched amount, so the adjusted total generation profile balances the grid. We denote the real-time adjustment by $\mathbf{g}\in\mathbb{R}^{N_{\mathrm{B}}}$. Mathematically, the two-stage network RLD problem takes the form of
@@
where $\bm{\alpha}\in\mathbb{R}^{N_{\mathrm{B}}}$ is the vector of linear generation cost coefficients for the day-ahead stage, and $Q(\mathbf{u},\mathbf{d})$ is the optimal second-stage cost given the day-ahead dispatch $\mathbf{u}$ and the realized net demand $\mathbf{d}$, defined as the optimal value of the real-time dispatch problem. In particular, as in [8], we use the following optimal DC (i.e., linearized AC) power flow formulation for the real-time dispatch:
@@
$-\mathbf{f}^{\mathrm{max}}\leq\mathbf{F}\bm{\theta}\leq\mathbf{f}^{\mathrm{max}}:\bm{\nu}_{-},\bm{\nu}_{+}.$ (2c)
Objective (2a) models the cost of generation adjustments or voluntary load shedding (like demand response) in real time, where $(\mathbf{g})_{+}:=\max\{\mathbf{g},\mathbf{0}\}$ and $\bm{\beta}\in\mathbb{R}^{N_{\mathrm{B}}}$ represents the per unit cost. Specifically, a positive component $g_{n}$ in $\mathbf{g}$ corresponds to two possible scenarios: i) Generator $n$ increases its output from $u_{n}$ to $u_{n}+g_{n}$, incurring an additional generation cost of $\beta_{n}g_{n}$, or ii) power demand at bus $n$ is curtailed from $d_{n}$ to $d_{n}-g_{n}$, incurring a penalty of $\beta_{n}g_{n}$. Thus, the term $\bm{\beta}^{\intercal}(\mathbf{g})_{+}$ captures the extra cost associated with these real-time actions. Eq. (2b) represents the DC power flow equation, where matrix $\mathbf{B}\in\mathbb{R}^{{N_{\mathrm{B}}}\times{N_{\mathrm{B}}}}$ maps voltage phase angles $\bm{\theta}\in\mathbb{R}^{N_{\mathrm{B}}}$ into nodal active power injections. Eq. (2c) denotes the branch capacity constraints, where $\mathbf{F}\in\mathbb{R}^{{N_{\mathrm{L}}}\times{N_{\mathrm{B}}}}$ maps $\bm{\theta}$ to branch flows, $\mathbf{f}^{\mathrm{max}}$ is the corresponding maximum allowable values, and ${N_{\mathrm{L}}}$ is the branch number. Vectors $\bm{\lambda}\in\mathbb{R}^{N_{\mathrm{B}}}$, $\bm{\nu}_{-}\in\mathbb{R}^{N_{\mathrm{L}}}$ and $\bm{\nu}_{+}\in\mathbb{R}^{N_{\mathrm{L}}}$ are the dual variables of (2b)-(2c), respectively. We use the DC power flow formulation instead of the AC formulation because, in practical wholesale markets, day-ahead generation decisions in transmission networks are typically made based on the DC model [25]. Moreover, the accuracy of the DC formulation is generally acceptable in transmission network operations, as line reactances are relatively small compared to resistances [26].
Remark 1 (VOLL interpretation)
Alternatively, the objective (2a) can be interpreted as the value of lost load (VOLL), in which case $\bm{\beta}$ is the per unit cost that users are willing to pay to avoid a loss of load event (involuntary load shedding). As a result, $\mathbb{E}_{\mathbf{d}}[Q(\mathbf{u},\mathbf{d})]$ models the risk associated with potential loss of load events [^1] VOLL is among one of the widest used power system risk metrics [27]. A power system operator that utilizes the objective in (1) to make its decision is risk-averse in the sense of [28, Definition 6.C.1] as it is easy to show that $Q(\mathbf{u},\mathbb{E}\mathbf{d})\leq\mathbb{E}Q(\mathbf{u},\mathbf{d})$ for any $\mathbf{u}$., and (1) trades off day-ahead generation cost with real-time operational risk. Note the values of $\bm{\beta}$ here are much higher compared to the voluntary load shedding scenario. Furthermore, the power system operator can reduce the risk of loss of load event by choosing larger values for $\bm{\beta}$.
@@
II-B Data-driven limiting dispatch
@@
Ideally, we wish to find the optimal decision rule $\mathbf{h}^{\star}:\mathbb{R}^{p}\mapsto\mathbb{R}^{N_{\mathrm{B}}}$, that maps the day-ahead auxiliary information $\mathbf{x}$ to the optimal dispatch decision $\mathbf{u}^{\star}$. Adapting terminologies from the statistical learning theory, we will also refer to such a decision rule as a hypothesis. Mathematically, supposing that there is an underlying true joint distribution of $(\mathbf{x},\mathbf{d})$, denoted by $\mathcal{Z}$, we can write the following problem to identify the optimal hypothesis as